Neptune Company produces toys and other items for use in beach and resort areas. A small, inflatable toy has come onto the market that the company is anxious to produce and sell. The new toy will sell for $2.70 per unit. Enough capacity exists in the company’s plant to produce 30,300 units of the toy each month. Variable expenses to manufacture and sell one unit would be $1.72, and fixed expenses associated with the toy would total $44,041 per month.

The company's Marketing Department predicts that demand for the new toy will exceed the 30,300 units that the company is able to produce. Additional manufacturing space can be rented from another company at a fixed expense of $2,202 per month. Variable expenses in the rented facility would total $1.89 per unit, due to somewhat less efficient operations than in the main plant.

Required:

1. What is the monthly break-even point for the new toy in unit sales and dollar sales.

2. How many units must be sold each month to attain a target profit of $9,963 per month?

3. If the sales manager receives a bonus of 15 cents for each unit sold in excess of the break-even point, how many units must be sold each month to attain a target profit that equals a 23% return on the monthly investment in fixed expenses?
Answer #1

Solution:-1

Contribution per unit for the first 30300 units= $2.7 - $1.72 = $0.98

Contribution margin in excess of 30300 units = $2.7 - $1.89 =$ 0.81

Total fixed costs= $44,041 + $2,202 = $46,243       

Recovery of fixed costs from first 30300 units = $0.98 * 30300 units = $29,694

Balance fixed costs to be recovered = $46,243 - $29694 = $16,549

Break even to recover balance costs = $16,549 / $0.81 = 20,431

Total break even units = 30300 units + 20,431units = 50,731 units

Break-even point in dollar sales= break even units * sale price per unit

Break-even point in dollar sales = 50,731 units * $2.7 = $136,974

Solution:-2

Required contribution margin= Total fixed cost +required target profit

Required contribution margin= $46,243+$9,963 = $56,206

Contribution required from new facility =$ 56,206 – $29,696 =$26,512     

Number of units required to break even from new facility = $26,512 /$ 0.81 = 32,731 units

Unit sales needed to attain target profit = 30300 + 32,731 = 63,031 units

Solution:-3

Contribution margin after bonus = 0.81% - 0.15% = 0.66

Desired profit = Total fixed expenses * rate of return = $46,243 * 23% = $10,636

No of units to break even to earn desired profit = $10,636 / 0.66 = 16,115 units

Unit sales needed to attain target profit = 50,731units + 16,115 units = 66,846 units